The radius of the tower is 15 meters. The height of the tower is 50 meters.
What is the volume of the water tower?

Answer by richwmiller(17219)

We do not know the shape of the water tower.
Is it a cylinder, a sphere or a cone?
Assuming that it is a cylinder
The volume is pi*r^2*h=pi*15^2*50=pi*225*50=pi*11250=35235 cubic meters